OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)? A)Q(1, 11) B)R(2, 4) C)S(4, -4) D)T(9, -2) PLZ HELP!!!

OpenStudy (anonymous):

@Breezyb123 @Squirrels @MrNood can you guys help me??

OpenStudy (amistre64):

subtract the center from all the points, then find the one that is most likely some variant of the 3,4,5 right triangle

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

its not a ....

OpenStudy (mrnood):

@Joscelyna Why would you trust someone who just writes "a" as a solution? - you have no way of knowing whether they have any competence at all. I have seem MANY such answers given here which are wrong - and the person who posted is at risk of putting THEIR scores in danger if they take the answer without checking. I will post below the method to find the answer - you should work it out for yourself - if you are not sure, then show your working and ask someone to check it...

OpenStudy (mrnood):

 you can see that the radius goes from the centre to any point on the circle and you can make a right angle triangle as shown in dotted lines. The length of the vertical side of the triangle is (y2-y1) The length of the horizontal side of the triangle is (x2-x1) In your question you know that the point in the centre is (6,1) i.e. x1=6 & y1=1 You are also given 4 points to try out (so for a x2=1 &y2=11) Now in any right angle trangle you know that a^2 + b^2 = c^2 where in this case c is the radius, which you are told is 5. SO for example in option a vertical side = y2-y1 = 11-1 = 10 horizontal side = x2-x1 = 1-6 = -5 so we need to check whether this makes a triangle with r = 5 10^2 + (-5)^2 = 100+25 = 125 r^2 = 5^2 = 25 so a^2 + b^2 is NOT equal to r^2 so the point in a is NOT on th ecircle. Do the same calculation for the other 3 options: in one of them you will fnd a^2 + b^2 = 25.